SHOW AND EXPLAIN ME THE ADDATION IN VECTORS

Grade:12

4 Answers

manu saxena
13 years ago
Two vectors can be added together to determine the result (or resultant).Vector addition involves finding vector components, adding them and finding the polar form of the resultant.
Neer Varshney
13 years ago

Its very easy respective elements are added:

like consider two matrices of 3*3. 

1 2 3              3 2 1

4 5 6    and   6 5 4

7 8 9             9 8 7

Then there add. will be:

4 4 4

10 10 10 

16 16 16. 

I hope u got it . Add first first , 2nd 2nd, 3rd 3rd and so on.

'

yash agarwal
13 years ago

Its very easy.

Suppose the vectors are given in their resolved forms like 2i+3j+5k and 1i+2j+4k

Just add the x,y and z components like this>>>

(2+1)i+(3+2)j+(5+4)k=3i+5j+9k this is your resultant vector.

its magnitude is sq.root of 9+25+81.

Or u can add the vectors graphically.Just position them head to tail and the resutant

will be from the empty tail to empty head.

xyz xz
12 years ago

IF A = ai+bj+ck and B = xi+yj+zk  where i,j,k represent unit vectors along x, y and z direction respectively

then

(A+B) = (a+x)i + (b+y)j + (c+z)k

magnitube of (A+B), IA+BI = (IAI2 + IBI2 + 2ABcos theta)1/2

 

i hope that u got a clear idea about addition of vectors...........Smile